# radarODE: An ODE-Embedded Deep Learning Model for Contactless ECG Reconstruction from Millimeter-Wave Radar

### v1

@@ IV-C3 Performance of radarODE
    Fine-Grained Cardiac Events Reconstruction: The evaluation of fine-grained cardiac features aims to analyze the temporal accuracy of the QRST peaks after filtering out the corrupt reconstructions. The P peak is not considered in this evaluation as also suggested in the benchmark paper [35], because the P peak is inconspicuous and even unable to be detected for some ground truth ECG pieces. The overall reconstructed ECG peak accuracy is shown in Figure 12a as CDF with the median/$90$-percentile absolute timing error for QRST peaks shown in Table IV. The results demonstrate that radarODE obtains a better peak accuracy than MMECG with the overall temporal error for QRST peaks improved by $5\%$, $23\%$, $37\%$, and $14\%$, respectively. The improvement is firstly attributed to the SST spectrogram with more evident patterns for reconstructing QRS-complex and T peak compared with using the time-domain signal as the input. Secondly, the ODE decoder itself cannot produce faithful reconstruction as shown in Figure 9a but could contribute to calibrating the peak positions according to $\eta$ and $\tau$, hence improving the overall peak accuracy. In addition, it is observed that the accuracy for the R peak is higher than the other peaks for both methods, because the deep neural network training relies on RMSE error and even a small deviation of the R peak with high amplitude causes large RMSE, hence the ‘learning’ of R peak becoming the most efficient way to reduce the RMSE loss.
    TABLE IV: Absolute Timing Error for Reconstructed ECG Peaks
@@
    Figure 12: Reconstructed ECG Peak Accuracy: (a) CDF of the overall temporal error for the reconstructed QRST peaks using radarODE and MMECG; (b) and (c) The peak error in terms of different physical statuses using radarODE and MMECG.
@@
    Figure 12b and 12c demonstrate the improvement of radarODE from the view of different physical statues, and the results coincide with many previous evaluations. Firstly, the radarODE performs better than MMECG for all statuses with R peak always achieving the best accuracy. Secondly, the stable physical statuses tend to yield accurate reconstruction with small variance, but the difference between statues is smaller than that of the morphological accuracy analysis, because the corrupt reconstructions have been filtered for peak accuracy evaluation.
V Conclusions
@@
Radar-based ECG reconstruction is a critical technique in many future applications such as smart home and in-cabin monitoring, but the existing methods all apply the pure data-driven approaches, lacking the theoretical support regarding the transformation between mechanical activities measured by radar and electrical activities described as ECG. This research aims to bridge the gap to realize the transformation from the mechanical domain to the electrical domain by proposing the signal model with fine-grained features considered and further designing the deep learning framework radarODE with morphological prior embedding in the form of ODEs. The radarODE framework is validated on the public dataset containing $4.5$ hours of radar measurements with corresponding ablation study, and the performance of the radarODE is also compared with the benchmark framework. The result evaluation shows that radarODE could achieve a better MDR, morphological accuracy and peak accuracy than the benchmark, proving the rationality of the proposed signal model and the effectiveness of the radarODE under various physical statuses. In the future, the accuracy of the PPI estimation needs to be improved by using deep-learning-based methods, and several data augmentation techniques might be applied to alleviate the data shortages, especially for PE status.

### v2

@@ V-C3 Fine-Grained Cardiac Events Reconstruction
The evaluation of fine-grained cardiac features aims to analyze the timing accuracy of the QRST peaks and the P peak is not considered in this evaluation as also suggested in the benchmark paper [28], because the P peak is inconspicuous and even unable to be detected for some ground truth ECG signal. The overall result is shown in Figure LABEL:sub@fig:peak_cdf as CDF with the median/$90$-percentile absolute timing error for QRST peaks shown in Table VII. The improvement owes to the use of SST spectrogram with more evident patterns for the prominent heart vibrations, and the ODE decoder contributes to calibrating the peak positions according to $\eta$ and $\tau$, hence improving the overall peak accuracy.
TABLE VII: Absolute Timing Error for Reconstructed ECG Peaks
@@
Figure 14: Evaluations of fine-grained cardiac events: (a) CDF of the overall timing error; (b) and (c) The timing error for different peaks under various physical statuses for radarODE and MMECG.
@@
Figure LABEL:sub@fig:peak_phy_ode and LABEL:sub@fig:peak_phy_mmecg demonstrate the improvement of radarODE in terms of different physical statuses, and the results coincide with many previous evaluations. Firstly, the radarODE outperforms MMECG for all statuses with the R peak always achieving the best accuracy. Secondly, stable physical statuses tend to yield accurate reconstruction with small variance, but the difference between statues is smaller than that of the morphological accuracy analysis, because the corrupt reconstructions have been filtered for peak accuracy evaluation.
V-C4 Summary of Long-term ECG Recovery Evaluation
The experimental results illustrated that the proposed radarODE could yield high-quality long-term ECG recovery even under extensive body movement noise, with better RMSE and PCC compared with the benchmark. In addition, the introduction of the SST spectrogram and the ODE decoder further improves the accuracy of the fine-grained cardiac features that are crucial to the potential applications in clinical diagnosis.
TABLE VIII: Complexity Analysis of Deep Learning Framework
Framework       Time Complexity Space Complexity Time/Epoch (min)
$\mathcal{O}(\cdot)$ per layer FLOPs (G) Params. (M)
MMECG [28] MHSA: $\mathcal{O}(CL^{2})$, Conv1d: $\mathcal{O}(LKC^{2})$ $0.59$ $0.67$ $3.25$
RSSRNet [29] MHSA: $\mathcal{O}(CL^{2})$, Conv2d:$\mathcal{O}(HLK^{2}C^{2})$ $1.06$ $4.51$ $4.02$
RadarNet [30] Conv1d: $\mathcal{O}(LKC^{2})$ $0.24$ $0.33$ $2.16$
radarODE Conv2d: $\mathcal{O}(HLK^{2}C^{2})$ $1.45$ $6.04$ $4.51$
MHSA: Multi-head self-attention, L: Length of input, H: Height of input, K: Kernel size, C: Channel
V-D Complexity Analysis and Comparison
Although the proposed radarODE performs better than other frameworks, it is necessary to analyze the complexity of the deep learning model for a fair comparison. In this subsection, the model complexity will be analyzed in terms of time and space complexity, and the training time for each epoch is also provided for an intuitive comparison as shown in Table VIII.
V-D1 Time Complexity
Time Complexity in deep learning typically refers to the computational cost during the model training and can be expressed in terms of floating-point operations (FLOPs) or Big O notation $\mathcal{O}(\cdot)$, with FLOPs measuring the floating-point calculations and $\mathcal{O}(\cdot)$ describing the asymptotic behavior of the model with respect to input size or hyperparameters [51, 50].
Table VIII first shows $\mathcal{O}(\cdot)$ for the commonly used structures (layers) (i.e., multi-head self-attention (MHSA), Conv1d and Conv2d) for ECG recovery [51, 50]. In theory, Conv1d used by MMECG and RadarNet has a low time complexity with 1-D input and kernel size, while Conv2d in radarODE and RSSRNet significantly increases the complexity because the model needs to deal with extra information from the frequency domain. However, the comparison of $\mathcal{O}(\cdot)$ between different models can be trivial if multiple structures are used.
In Table VIII, FLOPs are also provided to directly show the calculations required for one forward propagation. The results indicate that radarODE and RSSRNet perform more calculations compared with others, because radarODE and RSSRNet are based on the spectrogram inputs, while MMECG and RadarNet are based on 1D radar signals inputs. In addition, RSSRNet adopts single-channel input while radarODE owns a large backbone to process $50$ channels as input, further increasing the time complexity.
V-D2 Space Complexity
Space Complexity refers to the memory required to store weights and biases [4, 51], and can be easily revealed by counting the number of parameters in the model as shown in Table VIII. Similar to time complexity, RadarNet and MMECG have fewer parameters than RSSRNet and radarODE because of the 1-D input. The large parameter space of radarODE mainly comes from the powerful backbone with the ability to process multi-channel spectrogram inputs.
V-D3 Training Time per Epoch
According to the complexity analysis above, the proposed radarODE has larger FLOPs and should spend more time for each epoch. However, the differences in training time per epoch are smaller than other metrics, because the training sets of other frameworks are formed based on arbitrary radar/ECG segments with a step length of $0.15$s, while the radarODE is based on the segment in each single cardiac cycle. Therefore, the number of segments in the training set of other frameworks is around $48$k and is much larger than that of radarODE with $19$k training pairs. The discrepancy in the scale of training dataset indicates that the proposed framework based on single cardiac cycles could alleviate the introduction of redundant information, because the segments based on step length may cause homogeneous training samples without any contribution to enhance the diversity of the dataset, but too many similar samples may increase the risk of overfitting and degrade the generalization ability of the deep learning model [53].
V-E Discussions and Future work
The proposed radarODE framework has shown outstanding performance compared with the previous work to generate faithful ECG signals under noisy scenarios, while the potential limitation will be discussed in this part to encourage future improvements in radar-based ECG recovery for real-life situations and applications.
V-E1 Long-range ECG Monitoring
The direct impact brought by long-range monitoring is to reduce the SNR of the received radar signal according to the link budget analysis [52]. In addition, the cardiac location requires to be pre-identified to perform the accurate beamforming, because the current dataset assumes the majority of range-bins contain cardiac-related signals [28].
V-E2 Complex Monitoring Scenarios
Various new noises might be introduced and need to be eliminated, such as radar self-vibration introduced from car vibrations or hand-held radar [15], mutual-radar interference for the future smart home with multiple electromagnetic devices [16] and signal attenuation caused by human tissues for the monitoring of people with random body orientations [52].
V-E3 Evaluation on the Dataset for Patients
An important future application of radar-based ECG recovery is for clinical monitoring and diagnosis, while the ECG waveform for patients (e.g., arrhythmia) might be quite different, requiring massive new data for training. Some recent research has shown the feasibility of recovering abnormal ECG from radar signal [32], but more studies are required to investigate transfer learning or data augmentation techniques due to the scarcity of patient data. In addition, it is hard to preserve the noise-robustness for the ECG monitoring of patients, because the ODE model in this work is not designed for abnormal ECG patterns.
VI Conclusions
@@
Radar-based ECG reconstruction is highly reliant on purely data-driven approaches and lacks theoretical support regarding the transformation between mechanical activities measured by radar and electrical activities described as ECG. This research aims to bridge the gap to realize the transformation from the mechanical domain to the electrical domain by proposing the signal model with fine-grained features considered and further designing a deep learning framework radarODE with morphological prior embedding as ODEs. The radarODE framework is validated on the public dataset containing $4.5$ hours of radar measurements with corresponding ablation study and comparisons with the benchmark. The experimental result shows that radarODE could achieve a better MDR, morphological accuracy and peak accuracy than the benchmark, proving the rationality of the proposed signal model and the effectiveness of the radarODE under various physical statuses and random body movement. In the future, the complexity of the deep learning model needs to be reduced by squeezing the input size, and several data augmentation techniques might be applied to alleviate the data shortages, especially for patients.